# Feature Representation Transferring to Lightweight Models via Perception Coherence

### v1

@@ 5 Discussion and Conclusion
To sum up, we proposed an original method for deep representation transferring between models having arbitrary feature dimensions. This is achieved based on a novel concept named perception coherence. We propose a probabilistic framework to better understand the proposed notion. The theoretical results show that in the ideal scenario, the student model perceives the inputs in the same way as the teacher model. This provides a new way to theoretically understand the KD process. The experiments show promising results for lightweight KD. Moreover, our method is built generically for any dissimilarity metric for each of the student and teacher models. Hence, depending on application, it would be interesting to investigate task-specific metrics for a best KD, for example, in the multi-domain or multi-modal scenarios. We hope our work paves the way for future advancements in KD, especially based on topological deep learning frameworks.
Broader impact statement
Our work focuses on developing methods to distill knowledge from a neural networks to smaller, more efficient ones. By reducing the computational and memory requirements while maintaining good performance, these techniques have the potential to democratize access to advanced AI systems, making them more accessible across various applications. This could lead to advancements in areas such as on-device AI, or quick personalized services, through improved efficiency and effectiveness of AI technologies.
Appendix A Proof of Proposition 1
Proof.
By definition, we have that $F_{1}(x,X)=\mathbb{P}_{X^{\prime}}\left(d_{1}(x,X^{\prime})\leq d_{1}(x,X)\right)$ and $F_{2}(x,X)=\mathbb{P}_{X^{\prime}}\left(d_{2}(x,X^{\prime})\leq d_{2}(x,X)\right)$. As $f_{2}$ is absolutely perception coherent with $f_{1}$ at $x$, we obtain that $d_{1}(x,X^{\prime})\leq d_{1}(x,X)\Leftrightarrow d_{2}(x,X^{\prime})\leq d_{2}(x,X)$. Thus, $\mathbb{P}_{X^{\prime}}\left(d_{1}(x,X^{\prime})\leq d_{1}(x,X)\right)=\mathbb{P}_{X^{\prime}}\left(d_{2}(x,X^{\prime})\leq d_{2}(x,X)\right)$, which leads to the result. ∎
Appendix B Proof of Theorems 3.1 and 3.2
@@ Appendix A Proof of Proposition 1
Theorem B.1 (Theorem 3.1 in the main text).
@@
$\mathbb{P}_{X,X^{\prime}}\left(\left|F_{2}(x,X^{\prime})-F_{2}(x,X)\right|\geq\varepsilon_{2}\ \Bigm{|}\ \left|F_{1}(x,X^{\prime})-F_{1}(x,X)\right|\leq\varepsilon_{1}\right)\leq\frac{C_{\varepsilon_{1}}(1-\alpha)}{\varepsilon_{2}-\varepsilon_{1}}\ ,$ (B.1)
@@
Using Lemma 1, we have
@@
This completes the proof. ∎
@@
Theorem B.2 (Theorem 3.2 in the main text).
@@
$\mathbb{P}_{X,X^{\prime}}\left(F_{2}(x,X^{\prime})-F_{2}(x,X)\geq\varepsilon\ \big{|}\ F_{1}(x,X^{\prime})\leq F_{1}(x,X)\right)\leq C(1-\alpha)/\varepsilon\ ,$ (B.2)
@@
Using Lemma 1, we have
@@
This leads to the result of Theorem 3.2, which completes the proof. ∎
B.3 Lemma 1
Lemma 1.
@@ B.3 Lemma 1
$\leq\frac{2(1-\alpha)}{\varepsilon\mathbb{P}_{X,X^{\prime}}\left(A(X,X^{\prime})\right)}\ .$ (B.3)
Proof.
@@
Using Assumption 1, we have
@@
We note that $\left(\ B(X,X^{\prime})\Bigm{|}A(X,X^{\prime})\right)+\left(\ \overline{B(X,X^{\prime})}\Bigm{|}A(X,X^{\prime})\right)=1$. Using Assumption 1, we have that $\left(B(X,X^{\prime})\Bigm{|}A(X,X^{\prime})\right)+\left(B(X,X^{\prime})\Bigm{|}A(X,X^{\prime})\right)=1$. As $A(X,X^{\prime})$ is symmetric in $X$ and $X^{\prime}$, we obtain that $2\left(B(X,X^{\prime})\Bigm{|}A(X,X^{\prime})\right)=1$. Hence, $\left(B(X,X^{\prime})\Bigm{|}A(X,X^{\prime})\right)=1/2$. Therefore,
@@
[…]
@@
Note that by using Assumption 1 once again, we have
@@
This completes the proof. ∎
@@ Appendix C Proof of the theorem on global coherence
Theorem C.1 (Theorem 3.3 in the main text).
@@
Applying Lemma 2, we have that
@@
Applying Lemma 2 leads to
@@
where $C=\frac{4}{\mathbb{P}_{X_{1},X_{2},X_{3}}(A(X_{1},X_{2},X_{3}))}$. This completes the proof. ∎
C.2 Lemma 2
Lemma 2.
@@ C.2 Lemma 2
$\leq\frac{2(1-\alpha)}{\varepsilon\mathbb{P}_{X_{1},X_{2},X_{3}}\left(A(X_{1},X_{2},X_{3})\right)}\ .$ (C.1)
Proof.
@@
Using Assumption 1, we have
@@
×$\times\mathbb{P}_{X_{1},X_{2},X_{3}}\left(B(X_{1},X_{2},X_{3})\Bigm{|}A(X_{1},X_{2},X_{3})\right)$
@@
[…]
@@
This completes the proof. ∎
@@
Theorem D.1 (Theorem 3.4 in the main text).

### v2

@@ 6 Discussion and Conclusion
In summary, this paper introduced an original method for representation transfer between models with arbitrary feature dimensions, based on the novel concept of perception coherence. The probabilistic formulation provides new theoretical insights into the knowledge distillation process, showing that in the ideal case the student perceives the input space in the same way as the teacher. Experiments confirm that the method enables efficient lightweight transfer while maintaining competitive performance. Importantly, the proposed framework is generic and can be applied with different dissimilarity metrics for the teacher and student models. This flexibility suggests promising applications in task-specific contexts, such as multi-domain or multi-modal transfer. We believe that this work paves the way for future advancements in knowledge distillation, especially those grounded in topological perspectives for deep learning.
Appendix A Convergence of the global perception coherence level estimated by mini-batch
We provide here a full proof of Theorem 4.1.
Proof A.1
Step 1: Decompose error
We note that
$\widehat{\mathrm{DC}}_{B}-\mathrm{DC}$ $=\frac{1}{B^{2}}\sum_{i,j}\left(\left|\widehat{F}_{1,B}(X_{i},X_{j})-\widehat{F}_{2,B}(X_{i},X_{j})\right|-\left|F_{1}(X_{i},X_{j})-F_{2}(X_{i},X_{j})\right|\right)$
$\quad+\frac{1}{B^{2}}\sum_{i,j}\left|F_{1}(X_{i},X_{j})-F_{2}(X_{i},X_{j})\right|-\mathrm{DC}\ .$
Thus,
$\left|\widehat{\mathrm{DC}}_{B}-\mathrm{DC}\right|$ $\leq\left|\frac{1}{B^{2}}\sum_{i,j}\left(\left|\widehat{F}_{1,B}(X_{i},X_{j})-\widehat{F}_{2,B}(X_{i},X_{j})\right|-\left|F_{1}(X_{i},X_{j})-F_{2}(X_{i},X_{j})\right|\right)\right|$
$\quad+\left|\frac{1}{B^{2}}\sum_{i,j}\left|F_{1}(X_{i},X_{j})-F_{2}(X_{i},X_{j})\right|-\mathrm{DC}\right|.$
We now take expectation on both sides:
$\mathbb{E}\left|\widehat{\mathrm{DC}}_{B}-\mathrm{DC}\right|\leq E_{1}+E_{2},$
where
$E_{1}$ $:=\mathbb{E}\left[\left|\frac{1}{B^{2}}\sum_{i,j}\left(\left|\widehat{F}_{1,B}(X_{i},X_{j})-\widehat{F}_{2,B}(X_{i},X_{j})\right|-\left|F_{1}(X_{i},X_{j})-F_{2}(X_{i},X_{j})\right|\right)\right|\right],$
$E_{2}$ $:=\mathbb{E}\left[\left|\frac{1}{B^{2}}\sum_{i,j}\left|F_{1}(X_{i},X_{j})-F_{2}(X_{i},X_{j})\right|-\mathrm{DC}\right|\right].$
Step 2: Bound $E_{1}$
We use the reverse triangle inequality: for any real numbers $a,b$,
$\left||a|-|b|\right|\leq|a-b|.$
Applying this, we obtain:
$E_{1}$ $=\mathbb{E}\left[\left|\frac{1}{B^{2}}\sum_{i,j=1}^{B}\left(\left|\widehat{F}_{1,B}(X_{i},X_{j})-\widehat{F}_{2,B}(X_{i},X_{j})\right|-\left|F_{1}(X_{i},X_{j})-F_{2}(X_{i},X_{j})\right|\right)\right|\right]$
$\leq\mathbb{E}\left[\frac{1}{B^{2}}\sum_{i,j=1}^{B}\left|\left(\widehat{F}_{1,B}(X_{i},X_{j})-\widehat{F}_{2,B}(X_{i},X_{j})\right)-\left(F_{1}(X_{i},X_{j})-F_{2}(X_{i},X_{j})\right)\right|\right]$
$=\mathbb{E}\left[\frac{1}{B^{2}}\sum_{i,j=1}^{B}\left|\left(\widehat{F}_{1,B}(X_{i},X_{j})-F_{1}(X_{i},X_{j})\right)-\left(\widehat{F}_{2,B}(X_{i},X_{j})-F_{2}(X_{i},X_{j})\right)\right|\right]$
$\leq\mathbb{E}\left[\frac{1}{B^{2}}\sum_{i,j=1}^{B}\left(\left|\widehat{F}_{1,B}(X_{i},X_{j})-F_{1}(X_{i},X_{j})\right|+\left|\widehat{F}_{2,B}(X_{i},X_{j})-F_{2}(X_{i},X_{j})\right|\right)\right]$
$=\frac{1}{B^{2}}\sum_{i,j=1}^{B}\left(\mathbb{E}\left[\left|\widehat{F}_{1,B}(X_{i},X_{j})-F_{1}(X_{i},X_{j})\right|\right]+\mathbb{E}\left[\left|\widehat{F}_{2,B}(X_{i},X_{j})-F_{2}(X_{i},X_{j})\right|\right]\right)\ .$
Control empirical CDF deviations. For fixed $i\in\{1,2\}$ and fixed $x,x^{\prime}\in\mathcal{X}$, recall that the empirical CDF is defined as:
$\widehat{F}_{i,B}(x,x^{\prime}):=\frac{1}{B}\sum_{k=1}^{B}\mathds{1}\left\{d_{i}(x,X_{k})\leq d_{i}(x,x^{\prime})\right\},$
where $X_{1},\dots,X_{B}\sim\mathcal{D}$ i.i.d.
Each summand is a Bernoulli random variable with mean:
$\mathbb{E}\left[\mathds{1}\left\{d_{i}(x,X_{k})\leq d_{i}(x,x^{\prime})\right\}\right]=\mathbb{P}(d_{i}(x,X_{k})\leq d_{i}(x,x^{\prime}))=F_{i}(x,x^{\prime}).$
Hence, $\widehat{F}_{i,B}(x,x^{\prime})$ is the average of $B$ i.i.d. Bernoulli random variables with parameter $F_{i}(x,x^{\prime})$, and has variance:
$\operatorname{Var}\left(\widehat{F}_{i,B}(x,x^{\prime})\right)=\frac{1}{B}F_{i}(x,x^{\prime})(1-F_{i}(x,x^{\prime}))\leq\frac{1}{4B}.$
By Jensen’s inequality with function $t\mapsto t^{2}$ convex, we have:
$\mathbb{E}\left|\widehat{F}_{i,B}(x,x^{\prime})-F_{i}(x,x^{\prime})\right|\leq\sqrt{\operatorname{Var}(\widehat{F}_{i,B}(x,x^{\prime}))}\leq\frac{1}{2\sqrt{B}}.$
As this hold for any $x,x^{\prime}$, for each fixed $i\in\{1,2\}$, we therefore obtain
$E_{1}\leq\frac{1}{B^{2}}\sum_{i,j}\left(\frac{1}{2\sqrt{B}}+\frac{1}{2\sqrt{B}}\right)=\frac{B^{2}}{B^{2}}\cdot\frac{1}{\sqrt{B}}=\frac{1}{\sqrt{B}}.$
Step 4: Bound $E_{2}$
Using Lemma A.1, there exists a constant $C^{\prime}$ (independent of $B$) such that
$E_{2}\leq\frac{C^{\prime}}{\sqrt{B}}.$
Final bound:
$\mathbb{E}\left[\left|\frac{1}{B^{2}}\sum_{i,j=1}^{B}\left|F_{1}(X_{i},X_{j})-F_{2}(X_{i},X_{j})\right|-\mathrm{DC}\right|\right]\leq E_{1}+E_{2}\leq\frac{1+C^{\prime}}{\sqrt{B}}=\frac{C}{\sqrt{B}}\ .$
Remark A.1
No assumptions about ties or continuity of $d_{i}$ are needed for this expectation convergence.
A.1 Technical lemma
Lemma A.1
Let $X_{1},\dots,X_{B}$ be i.i.d. draws from a distribution $\mathcal{D}$ on a measurable space $\mathcal{X}$. Let $h:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$ be a measurable kernel (not assumed symmetric). Define $\theta:=\mathbb{E}\big[h(X_{1},X_{2})\big]$. Assume $h$ is uniformly bounded: $\|h\|_{\infty}:=\sup_{x,y}|h(x,y)|=:M<\infty$. We define
$V_{B}\;:=\;\frac{1}{B^{2}}\sum_{i=1}^{B}\sum_{j=1}^{B}h(X_{i},X_{j}).$
Then,
$\mathbb{E}\big[\,|V_{B}-\theta|\,\big]\leq\frac{2(\sqrt{2}+6\sqrt{3})M}{\sqrt{B}}.$
In particular, $\mathbb{E}|V_{B}-\theta|=O(B^{-1/2})$.
Proof A.2
The proof follows the standard Hoeffding (projection) decomposition and a counting argument for the variance of the degenerate component; no symmetry of $h$ is required because we sum over all ordered pairs $(i,j)$.
1. Hoeffding decomposition (ordered pairs).
Let $X^{\prime}$ denote an independent copy of $X_{1}$. Define the first-order projection
$h_{1}(x):=\mathbb{E}\big[h(x,X^{\prime})\big]-\theta,$
and the second-order (degenerate) remainder
$h_{2}(x,y):=h(x,y)-h_{1}(x)-\tilde{h}_{1}(y)-\theta,$
where for notational symmetry we also set
$\tilde{h}_{1}(y):=\mathbb{E}\big[h(X^{\prime},y)\big]-\theta.$
By construction
$\mathbb{E}[h_{1}(X)]=\mathbb{E}[\tilde{h}_{1}(X)]=0,\qquad\mathbb{E}[h_{2}(X,y)]=\mathbb{E}[h_{2}(x,X)]=0\quad\text{for all }x,y.$
The kernel decomposes as
$h(x,y)=\theta+h_{1}(x)+\tilde{h}_{1}(y)+h_{2}(x,y).$
2. Decomposition of the V-statistic.
Plugging this into $V_{B}$ gives
$\begin{split}V_{B}-\theta&=\frac{1}{B^{2}}\sum_{i,j}\big(h_{1}(X_{i})+\tilde{h}_{1}(X_{j})+h_{2}(X_{i},X_{j})\big)\\ &=\frac{1}{B^{2}}\Big(B\sum_{i=1}^{B}h_{1}(X_{i})+B\sum_{j=1}^{B}\tilde{h}_{1}(X_{j})\Big)+\frac{1}{B^{2}}\sum_{i,j}h_{2}(X_{i},X_{j})\\ &=\frac{1}{B}\sum_{i=1}^{B}h_{1}(X_{i})+\frac{1}{B}\sum_{j=1}^{B}\tilde{h}_{1}(X_{j})+R_{B},\end{split}$
where we denote
$R_{B}:=\frac{1}{B^{2}}\sum_{i,j=1}^{B}h_{2}(X_{i},X_{j}).$
Define the linear part $L_{B}:=\frac{1}{B}\sum_{i}h_{1}(X_{i})+\frac{1}{B}\sum_{j}\tilde{h}_{1}(X_{j})$, so $V_{B}-\theta=L_{B}+R_{B}$.
3. Bound the linear part $L_{B}$.
The two sample sums are independent and mean zero; hence
$\operatorname{Var}(L_{B})=\frac{1}{B^{2}}\cdot B\operatorname{Var}(h_{1}(X))+\frac{1}{B^{2}}\cdot B\operatorname{Var}(\tilde{h}_{1}(X))=\frac{1}{B}\big(\operatorname{Var}(h_{1}(X))+\operatorname{Var}(\tilde{h}_{1}(X))\big).$
By Jensen,
$\mathbb{E}[|L_{B}|]\leq\sqrt{\operatorname{Var}(L_{B})}=\frac{1}{\sqrt{B}}\sqrt{\operatorname{Var}(h_{1}(X))+\operatorname{Var}(\tilde{h}_{1}(X))}.$
It is easy to see that $|h_{1}|\leq 2M$. Hence, $\operatorname{Var}(h_{1}(X))\leq 4M^{2}$. Similarly, $\operatorname{Var}(\tilde{h}_{1}(X))\leq 4M^{2}$. Therefore
$\mathbb{E}[|L_{B}|]\leq\frac{2\sqrt{2}M}{\sqrt{B}}.$
4. Bound the degenerate part $R_{B}$.
We bound $\mathbb{E}[|R_{B}|]$ via its second moment:
$\mathbb{E}[|R_{B}|]\leq\sqrt{\mathbb{E}[R_{B}^{2}]}=\sqrt{\operatorname{Var}(R_{B})}.$
Compute
$\operatorname{Var}(R_{B})=\frac{1}{B^{4}}\sum_{i,j,k,\ell=1}^{B}\operatorname{Cov}\!\big(h_{2}(X_{i},X_{j}),\,h_{2}(X_{k},X_{\ell})\big).$
Independence implies that if the index sets $\{i,j\}$ and $\{k,\ell\}$ are disjoint (i.e. all four indices are distinct), then the covariance is zero. Thus only quadruples with at least one shared index contribute. We count these quadruples.
The number of ordered quadruples with all four indices distinct equals
$B(B-1)(B-2)(B-3)=B^{4}-6B^{3}+11B^{2}-6B,$
so the number of quadruples with at least one common index is
$B^{4}-\big(B^{4}-6B^{3}+11B^{2}-6B\big)=6B^{3}-11B^{2}+6B\leq 12B^{3}.$
Notice that $\mathbb{E}\left[h_{2}(X_{i},X_{j})\right]=0$. Therefore,
$\big|\operatorname{Cov}\left(h_{2}(X_{i},X_{j})h_{2}(X_{k},X_{\ell})\right)\big|=\big|\mathbb{E}\left[h_{2}(X_{i},X_{j})\cdot h_{2}(X_{k},X_{\ell})\right]\big|\leq\sup|h_{2}|^{2}\ .$
Moreover, $|h_{2}|\leq|h|+|h_{1}|+|\tilde{h}_{1}|+|\theta|$. It is easy to see that $|\theta|\leq M$, $|h_{1}|\leq 2M$ and $|\tilde{h}_{1}|\leq 2M$. That is, $|h_{2}|\leq 6M$. Therefore,
$\big|\operatorname{Cov}\left(h_{2}(X_{i},X_{j})h_{2}(X_{k},X_{\ell})\right)\big|\leq 36M^{2}\ .$
Consequently,
$\operatorname{Var}(R_{B})\leq\frac{12B^{3}}{B^{4}}\cdot 36M^{2}=\frac{(12\cdot 36)M^{2}}{B}.$
Therefore
$\mathbb{E}[|R_{B}|]\leq\sqrt{\operatorname{Var}(R_{B})}\leq\frac{12\sqrt{3}\,M}{\sqrt{B}}.$
5. Combine the two bounds.
Using the triangle inequality,
$\mathbb{E}\big[\,|V_{B}-\theta|\,\big]\leq\mathbb{E}[|L_{B}|]+\mathbb{E}[|R_{B}|].$
Plugging (1) and (2) yields
$\mathbb{E}\big[\,|V_{B}-\theta|\,\big]\leq\frac{M}{\sqrt{B}}\Big(2\sqrt{2}+12\sqrt{3}\Big)=\frac{2(\sqrt{2}+6\sqrt{3})M}{\sqrt{B}}.$
Appendix B Proof of Theorems 4.2 and 4.3
@@ A.1 Technical lemma
Theorem B.1 (Theorem 4.2 in the main text)
@@
$\mathbb{P}_{X,X^{\prime}}\left(\left|F_{2}(x,X^{\prime})-F_{2}(x,X)\right|\geq\varepsilon_{2}\ \Bigm|\ \left|F_{1}(x,X^{\prime})-F_{1}(x,X)\right|\leq\varepsilon_{1}\right)\leq\frac{C_{\varepsilon_{1}}(1-\alpha)}{\varepsilon_{2}-\varepsilon_{1}}\ ,$ (8)
@@
Using Lemma B.1, we have
@@
This completes the proof.
@@
Theorem B.2 (Theorem 4.3 in the main text)
@@
$\mathbb{P}_{X,X^{\prime}}\left(F_{2}(x,X^{\prime})-F_{2}(x,X)\geq\varepsilon\ \big|\ F_{1}(x,X^{\prime})\leq F_{1}(x,X)\right)\leq C(1-\alpha)/\varepsilon\ ,$ (9)
@@
Using Lemma B.1, we have
@@
This leads to the result of Theorem 4.3, which completes the proof.
B.3 Lemma B.1
Lemma B.1
@@
$\leq\frac{2(1-\alpha)}{\varepsilon\mathbb{P}_{X,X^{\prime}}\left(A(X,X^{\prime})\right)}\ .$ (10)
Proof B.3
@@
Using Assumption 4.2, we have
@@
We note that $\left(\ B(X,X^{\prime})\Bigm|A(X,X^{\prime})\right)+\left(\ \overline{B(X,X^{\prime})}\Bigm|A(X,X^{\prime})\right)=1$. Using Assumption 4.2, we have that $\left(B(X,X^{\prime})\Bigm|A(X,X^{\prime})\right)+\left(B(X,X^{\prime})\Bigm|A(X,X^{\prime})\right)=1$. As $A(X,X^{\prime})$ is symmetric in $X$ and $X^{\prime}$, we obtain that $2\left(B(X,X^{\prime})\Bigm|A(X,X^{\prime})\right)=1$. Hence, $\left(B(X,X^{\prime})\Bigm|A(X,X^{\prime})\right)=1/2$. Therefore,
@@
$\leq\frac{\mathbb{E}_{X}\left[\left|F_{1}(x,X)-F_{2}(x,X)\right|\right]}{\varepsilon\mathbb{P}_{X,X^{\prime}}\left(A(X,X^{\prime})\bigcap B(X,X^{\prime})\right)}\;\text{(Markov's inequality)}$
@@
Note that by using Assumption 4.2 once again, we have
@@
This completes the proof.
@@ Appendix C Proof of the theorem on global coherence
Theorem C.1 (Theorem 4.4 in the main text)
@@
Applying Lemma C.1, we have that
@@
Applying Lemma C.1 leads to
@@
where $C=\frac{4}{\mathbb{P}_{X_{1},X_{2},X_{3}}(A(X_{1},X_{2},X_{3}))}$. This completes the proof.
C.2 Lemma C.1
Lemma C.1
@@
$\leq\frac{2(1-\alpha)}{\varepsilon\mathbb{P}_{X_{1},X_{2},X_{3}}\left(A(X_{1},X_{2},X_{3})\right)}\ .$ (11)
Proof C.2
@@
Using Assumption 4.2, we have
@@
$\times\mathbb{P}_{X_{1},X_{2},X_{3}}\left(B(X_{1},X_{2},X_{3})\Bigm|A(X_{1},X_{2},X_{3})\right)$
@@
$\leq\frac{\mathbb{E}_{X_{1}}\left[\frac{\mathbb{E}_{X_{2}}\left[\left|F_{1}(X_{1},X_{2})-F_{2}(X_{1},X_{2})\right|\right]}{\varepsilon}\right]}{\mathbb{P}_{X_{1},X_{2},X_{3}}\left(A(X_{1},X_{2},X_{3})\bigcap B(X_{1},X_{2},X_{3})\right)}\;\text{(Markov's inequality)}$
@@
This completes the proof.
@@
Theorem D.1 (Theorem 4.5 in the main text)